OpenStudy (safari321):

The width of a soccer field must be between 55 yd and 80 yd. a. what compound inequality represents the width of a soccer field?

OpenStudy (cruffo):

Let x = width of soccer field \[\large 55 \leq x \leq 80\]

OpenStudy (safari321):

I don't get it

OpenStudy (anonymous):

look at the word in the sentence, they give you clues to which signs or inequalities you should use. The clue is must be between so it can range from anything equal to 55 and greater and anything equal to 80 or less. So the key words there are "greater and equal to" and "less than or equal to." The width of the soccer field (x) is greater than or equal to 55 as said before and it is less than or equal to 80, this is written as \large 55 \leq x \leq 80

OpenStudy (anonymous):

that didnt show up right

OpenStudy (anonymous):

55≤x≤80 because it greater than or equal to 55 but less than or equal to 80

OpenStudy (safari321):

A little late but thanks.............